OpenStudy (anonymous):

Solve the equation 2 over 3 k + 6 = -12 for k. Please show your work and solution in the space provided

OpenStudy (anonymous):

2 / (3k+6) = -12 2 / -12 = 3k +6 -1/6 = 3k + 6 -1/6 - 6 = 3k -6 1/6 = 3k k = ?

OpenStudy (anonymous):

Solve the equation Quantity 4 x plus 8 all over 2 = -10 for x. Please show your work and solution in the space provided